A lab technician has 40% alcohol solution and 60% alcohol solution. He need to make 100 ml. solution that is 48% alcohol. How much of each available solution should he use ?
Answer by mananth(16946)

percent ---------------- quantity
Solution I 40.00% ---------------- x ml
solution II 60.00% ------ 100 - x ml
Mixture 48.00% ---------------- 100
Total 100 ml
40.00% x + 60.00% ( 100 - x ) = 48.00% * 100
40 x + 60 ( 100 - x ) = 4800
40 x + 6000 - 60 x = 4800
40 x - 60 x = 4800 - -6000
-20 x = -1200
/ -20
x = 60 ml 40.00% Solution I
40 ml 60.00% solution II

m.ananth@hotmail.ca